OpenStudy (anonymous):

Create your own third degree polynomial that when divided by x + 2 has a remainder of –4

OpenStudy (anonymous):

yessir, lol

OpenStudy (saifoo.khan):

;D

OpenStudy (saifoo.khan):

6.02 i think?

OpenStudy (anonymous):

Need some help getting this done - was supposed to be turned in last week and shes giving me extra time

OpenStudy (anonymous):

6.04 :)

OpenStudy (jamesj):

If a polynomial p(x) when divided by x+2 has remainder -4, then that means we can write p(x) = (x+2)q(x) + (-4) for some other polynomial q(x) Hence to answer your question, choose any q(x) in such a way that p(x) is third order.

OpenStudy (anonymous):

(x+2)(x^2 - 1) = x^3 - x + 2x^2 - 2 lets add 4 to that: x^3 + 2x^2 - x + 2 ill fit the bill